Yazmin noticed her favorite workout tops hereon sale for 6.00 off the original price. She purchased 4 of the tops and spent a total of 51.80. Write and solve an equation to determine x, the original price.
Answer by ikleyn(52778)

.
Yazmin noticed her favorite workout tops hereon sale for 6.00 off the original price.
She purchased 4 of the tops and spent a total of 51.80.
Write and solve an equation to determine x, the original price.
~~~~~~~~~~~~~~~~~~~

Let x be the original price per one top.

Then the price after discount is (x-6), and the equation for the purchase is

    4(x-6) = 51.80.


To solve, first divide both sides by 4

    x - 6 = 51.80/4

    x - 6 = 12.95


and find x

    x = 12.95 + 6 = 18.95.


ANSWER.  The original price is/was 18.95.

Solved, with explanations.